A roller coaster, loaded with passengers, has a mass of 2 500 kg; the radius of curvature of the track at the bottom point of the dip is 24 m. If the vehicle has a speed of 18 m/s at this point, what force is exerted on the vehicle by the track? (g = 9.8 m/s2)

2.3 x 104 N

4.7 x 104 N

5.8 x 104 N

7.3 x 104 N
Answer #1

F=N = mv2/r +mg =2500*(18^2/24+9.8)=  58250 N

answer =5.8 x 104 N